Bung Tie is running along a straight line. Assume that he starts from rest from point A and accelerates uniformly for T s, before reaching a speed  of 3 m/s. He is able to maintain this speed for 5 s. After that, it takes him 6 s to decelerate uniformly to come to a stop at point B. 

i. Sketch a speed versus time graph based on the information given above.

 

ii.Find the value of T if the distance between A and B is 100 m.

 

iii. Determine the deceleration.

 


1
Expert's answer
2022-06-13T08:22:58-0400

(i)



(ii)


"d=d_1+d_2+d_3\\
=\frac{3}{2}T+3*5+\frac{3}{2}*6=100"

"T=51\:\rm s"

(iii)

"a=\frac{v_f-v_i}{t}=\frac{0-3}{6}=-0.5\:\rm m/s^2"